A satellite of mass M orbits the earth with period T. A second satellite of mass 2M

is at the same radial distance from the center of the earth as the first satellite. What is the

period of the second satellite?

A. 1

2T

B. T

C. 2T

D. 4T


1
Expert's answer
2021-11-21T17:25:32-0500

"G\frac{Mm}{r^2}=m\omega_1^2r"


"G\frac{M2m}{r^2}=2m\omega_2^2r" . So, we have


"\omega_1^2/\omega_2^2=1\to \omega_1=\omega_2\to T_1=T_2=T"



B. T . Answer